OpenStudy (darkbluechocobo):

The height h of an arrow in feet is modeled by h(t) = -16t2 + 63t + 4, where t is the time in seconds since the arrow was shot. How long is the arrow in the air?

OpenStudy (da_scienceman):

so if I were you, I will find dh/dt = maximum of the resulting curve. Such maximum occurs ath the highest point of the given expression, where the velocity is zero. Then I proceed by solving dh(t)/dt = 0. This probably given me the value of t at this point, right? Then I assume that the time taken to arrive at this point corresponds to the value obtained at the maximum!

OpenStudy (darkbluechocobo):

so do I first divide 63/-16?

OpenStudy (da_scienceman):

No its a calculus question. U find the derivative of H(t) with respect to t. This gives you the velocity. Did you guys study calculus?

OpenStudy (darkbluechocobo):

Not really. This is my first application question I have done for this section.

OpenStudy (darkbluechocobo):

So where would i start solving this. Because I am really confused

OpenStudy (da_scienceman):

So you never heard about the derivative or what?

OpenStudy (darkbluechocobo):

no

OpenStudy (da_scienceman):

ok lets see u might need to solve the h(t) = 0 and probably divide by 2 or something. This could give u the time it takes. Hang on a sec lemme read the question again.

OpenStudy (darkbluechocobo):

how would you solve that? /.\ h(t)=0/2?

OpenStudy (jdoe0001):

notice that the leading coefficient of the parabola equation is negative, thus is opening DOWNWARD, so 

OpenStudy (darkbluechocobo):

Yes because the arrow would be falling

OpenStudy (jdoe0001):

 those would be the x-intercepts, so set y = 0, and solve for "x", like for any quadratic equaation \(\bf h(t) = -16t2 + 63t + 4\implies 0 = -16t2 + 63t + 4\)